OpenStudy (anonymous):

classify the conic section and wrote its equation in standard form. x^2+y^2+8x+12y+3=0

OpenStudy (campbell_st):

this is a circle, you will need to complete the square for x and y so that you can get it in the general form \[(x - h)^2 + (y - k)^2 = r^2\] so rewrite the equation as \[x^2 + 8x + y^2 + 12y = -3\] now complete the squares by adding (8/2)^2 and (12/2)^2 to both sides of the equation,

OpenStudy (anonymous):

so (x-4)^2+(y-6)^2=49 then what

OpenStudy (campbell_st):

thats about it... you now know the centre (4, 6) and the radius is 7